# E4S: Fine-grained Face Swapping via Editing With Regional GAN Inversion

### v2

@@ 3.4 Training Objective
Different from the most current face swapping methods, our E4S takes reconstruction as the proxy task, making it easy to train. When training is finished, one can employ the texture encoder $F_{\phi}$ to generate per-region texture codes for any input face. Then, face swapping can be easily fulfilled as described in Sec. 3.1. We apply the commonly used loss functions in the GAN inversion methods, which are detailed in our Appendix, where the training objectives of the face re-coloring network $B_{\psi}$ and face inpainting network $P_{\tau}$ are also presented.
4 Experiment Setup
4.1 DataSets.
@@ 4 Experiment Setup
FFHQ [47] contains 70K high-quality at $1024^{2}$ resolution. It includes vastly variation in terms of age, ethnicity and image background, and also has better coverage of accessories such as eyeglasses, sunglasses, hats, etc. Since it does not provide the facial segmentation masks, we use a pre-trained face parser [48] to obtain the facial segmentation masks.
4.2 Implementation Details.
The proposed RGI is trained with PyTorch [53] with 8 NVIDIA Tesla A100 GPUs. The batch size is 2 at each GPU during training. The learning rate is initialized as $10^{-4}$ with the Adam [54] optimizer ($\beta_{1}=0.9$, $\beta_{2}=0.999$). We train the model for 200K and 300K iterations on CelebAMask-HQ and FFHQ datasets, respectively. Besides, The initial learning rate decays by the factor of $0.1$ at 100K and 150K iterations, separately. Here, all images are randomly flipped with a ratio of $0.5$. As for the adversarial training, we update the parameters of the discriminator once the generator gets updated 15 times. The reason here is that the pre-trained StyleGAN can well preserve the texture details of inner face components, while sometimes its performance is unstable on preserving the hair details. Hence, we fine-tune the first $K=13$ layers to improve the hair quality. As for the RGI-Optimization, we fix the learning rate as $1e^{-2}$. Empirically, approximately 50 steps of optimization would bring satisfying results.
@@
Configurations SSIMnormal-$\uparrow$ PSNRnormal-$\uparrow$ RMSEnormal-$\downarrow$ FIDnormal-$\downarrow$
@@
The Role of Re-enactment. We first study the roles of re-enactment step of E4S framework, which aims to drive the source to show a similar pose and expression as the target. However, aside from pose and expression, an ideal reenactment model should not affect other source attributes, such as identity. Specifically, we employ a pre-trained face reenactment model [17] before the shape and texture swapping procedure. To verify the necessity of the re-enactment step in E4S, we compare a standard E4S pipeline and the one without re-enactment. As shown in the 1st column of Fig. 7, the swapped result is not aligned with the target face when the reenactment is disabled.
Figure 8: Qualitative results of ablation study on face re-coloring (zooming in for more details).
Figure 9: Qualitative results of ablation study on face inpainting (zooming in for more details).
@@
Pre-trained vs fine-tuned StyleGAN. The pre-trained StyleGAN can be used for face swapping. However, we notice that the hair texture details cannot be always well preserved. For a more robust performance on hair, we fine-tune the first $K=13$ layers of the StyleGAN. In Table I, we use a configuration (A) to indicate freezing the parameters of the StyleGAN generator and only training the texture encoder $F_{\phi}$ and the subsequent MLPs in our RGI. Although better SSIM, PSNR, and RMSE can be achieved by (A), the FID is poor. The last column in Fig. 7 also supports this. As contrast, fine-tuning can improve hair quality while maintaining the texture of other inner facial components. The generated images look unrealistic and lack certain textures (especially in hair). The generation quality can also be proved by he higher FID.
The Role of face re-coloring network $B_{\psi}$. To evaluate the performance of the proposed face re-coloring network $B_{\psi}$, we provide the qualitative results in Fig. 8. The comparison results demonstrate that using re-coloring network can produce more natural and target-consistent lighting. Besides, although $B_{\psi}$ only generates $256^{2}$-resolution images, the final re-coloring results have the comparable image quality with the $1024^{2}$ sized inputs, with the help of face enhancing network [50] restoring the details and re-coloring mask preserving the high-frequency information.
The Role of face inpainting network $P_{\tau}$. We also study the role of face inpainting network $P_{\tau}$ in Fig. 9. This clearly shows that as post-processing, network $P_{\tau}$ can successfully and adaptively keep the face shape of the source, even though the face-shape difference between source and target is extremely large. Besides, the quantitative comparison between our standard E4S and the one without face re-coloring network $B_{\psi}$ or face inpainting network $P_{\tau}$ are demonstrated in Table II. We would elaborate on this in Sec. 6.2.
@@
Figure 11: Comparisons with StyleGAN-based face swapping methods.
6 Face Swapping Results.
@@
We first illustrate several face swapping examples synthesized by our E4S framework and several representative methods in Fig. 10. Our method can realize high-quality swapping at the resolution of 1024$\times$1024. Then, we compare our method with the previous StyleGAN-based face swapping works in Fig. 11. Next, we randomly sample 500 source-target pairs from the test set of CelebAMask-HQ to obtain the swapped results of different approaches. Besides, we provide video face swapping results in Appendix. Moreover, we develop a user-interface system to perform image and video face swapping, which are also detailed in Appendix.
6.1 Qualitative comparison.
@@
We provide the qualitative comparisons with state-of-the-art methods in Fig. 10, where our E4S achieves more realistic and high-fidelity swapped results than the others. Compared with Compared with FSGAN [15], our results are much sharper. The results of SimSwap [8] and Hififace [10] suffer from some artifacts and distortions (see the 2nd and 4th row). Although both our E4S and FaceShifter [9] can generate visually-satisfying results, ours shows better detailed textures.
We further compare the performance in more challenging cases where the occlusion exists in the source and target faces (see the last two rows in Fig. 10). It can be observed that our E4S can fill out the missing skin in the source face (see the last row), and preserve the target glasses (see the 3rd row). As contrast, FSGAN results are quite blurry although they design a dedicated inpainting sub-network for occlusion. FaceShifter proposes a refinement network to maintain the occlusion in the target image, which also may bring back some identity information of the target, making the identity of their swapped results confused.
We also compare with some StyleGAN-based approaches in Fig. 11. The results of MegaFS [1] look to be a mixture of the source and target, which are blurry and lack of textures. StyleFuison [2] shows a bit of over-smoothing (see the last row). HiRes [3] still suffer from some artifacts in $1024^{2}$ resolution. As contrast, our E4S can generate more realistic and high-quality face swapping results.
TABLE II: Quantitative comparison with the existing face swapping methods. we denote the best results with bold marks. $\dagger$: source-oriented method, $\ddagger$: target-oriented method, *: StyleGAN-based method.
Method ID retrievalnormal-$\uparrow$ Posenormal-$\downarrow$ Expr.normal-$\downarrow$
Top-1 Top-5
$\dagger$ FSGAN [15] 0.17 0.32 2.33 2.45
$\ddagger$ SimSwap [8] 0.12 0.32 2.89 2.84
$\ddagger$ FaceShifter [9] 0.06 0.26 1.73 2.35
$\ddagger$ HifiFace [10] 0.15 0.37 2.77 2.82
MegaFS [1] 0.29 0.45 3.03 3.05
StyleFusion [2] 0.35 0.18 5.37 2.94
$\ddagger$ * HiRes [3] 0.05 0.51 2.71 2.83
$\dagger$ * Our E4S w/o $B_{\psi}$ 0.40 0.57 3.29 3.07
$\dagger$ * Our E4S w/o $P_{\tau}$ 0.38 0.55 3.28 3.06
$\dagger$ * Our standard E4S 0.42 0.58 3.28 3.05
@@
Figure 12: Qualitative comparisons of our results with state-of-the-art face editing methods. Some modifications are made on the face contour, hair, nose, and eyebrows. Our method can produce more high-fidelity editing results while maintaining the details of other components and the overall identity information well.
6.2 Quantitative comparison.
@@
The quantitative comparison is listed in Table II. It considers two aspects: identity preservation from the source and the attribute preservation from the target. Specifically, inspired by FaceShifter, our experimental settings are as follows. For source identity preservation, we first extract the ID feature vectors of all the source faces and the swapped results via CosFace [55]. Then, we conduct face retrieval by searching for the most similar face from all the source images for each swapped face, where the similarity is computed as the cosine distance. We use Top-1 and Top-5 accuracy are the evaluation metrics. For the target attribute preservation, we estimate the pose and expression by using HopeNet [56] and a 3D face reconstruction model [14], respectively. The $\ell_{2}$ distance of the pose and expression between each swapped face and its ground-truth target face is used as metrics.
@@
Table II shows our E4S achieves the best retrieval accuracy, which demonstrates the superiority of our method on identity preservation. As for the target attribute preservation, our E4S achieves comparable results in pose and expression. In general, the target-oriented methods usually keep more accurate pose and expression, since they start from the target face while our E4S is a source-oriented method, which generates the target attributes starting from the source. Nonetheless, one side effect of the target-oriented methods is that the injected identity information from the source is always limited (see FaceShifter in Fig. 10 and HiRes in Fig. 11). To sum up, there is a trade-off between identity and attribute preservation in source- and target-oriented methods. Note that our E4S has the potential to achieve better results by leveraging a more advanced reenactment method.
@@
Besides, we also evaluate how the proposed face re-coloring network $B_{\psi}$ and face inpainting network $P_{\tau}$ make a difference to the overall E4S framework in Table II. Although the face re-coloring network $B_{\psi}$ basically does not affect the performance on source ID retrieval and target pose and expression, it totally changes the visual effects of human eyes since the face skin tone is maintained from the target rather the source (see Fig. 8). Besides, Fig. 9 has demonstrated the proposed face inpainting network $P_{\tau}$ can successfully maintain the face shape, which cannot be solved well by mask-swapping only. The quantitative comparison in Table II demonstrates our face inpainting network $P_{\tau}$ can improve the performance of ID retrieval since accurately maintaining source face shape is important for source ID retrieval. Such a performance improvement is consistent with the results in Fig. 9.
7 Face Editing Results Via RGI.
@@
7.1 Qualitative comparison.
We show the visual comparison with the competing methods in Fig. 12. We make some modifications to the original facial mask, such as hair, eyebrows, and face contour. It shows that our approach produces more high-fidelity editing results, where the details of other components and the overall identity information are well kept.
TABLE III: Quantitative comparison for image reconstruction on CelebAMask-HQ [16] test set. The rows in gray indicate the reconstruction images are obtained via style code optimization.
method SSIMnormal-$\uparrow$ PSNRnormal-$\uparrow$ RMSEnormal-$\downarrow$ FIDnormal-$\downarrow$
@@
7.2 Quantitative comparison.
We make comparison between the competing methods and our RGI on the image reconstruction quality, where the Structural Similarity Index (SSIM) [58], Root-Mean-Squared-Error (RMSE), Peak Signal-to-Noise Ratio (PSNR), and Fréchet Inception Distance (FID) [59] are used as the metrics. The results are showed in Table III. We also compare E4S with SofGAN [46], which is a StyleGAN-based method and takes style code optimization for the reconstruction. For a fair comparison, we use RGI-Optimization. Table III shows that our method always achieves the best performance on all metrics, indicating the superiority of our method. SEAN [45] sometimes produces artifacts on hair regions while our RGI can achieve more high-fidelity reconstructions with better identity, texture, and illumination. Besides, our RGI-Optimization can preserve the facial details better (e.g., the curly degree of hair, the thickness of the beard, dimples, and background).
8 Conclusion
In this paper, we present a novel framework E4S for face swapping, which considers conducting face swapping from the perspective of fine-grained face editing, i.e., “editing for swapping”. Our E4S proposes to explicitly disentangle the shape and texture of each facial component, thus the face swapping task can be reformulated as a simplified problem of texture and shape swapping. To achieve such disentanglement as well as high resolution and high fidelity, we propose a novel Regional GAN inversion (RGI) method. Specifically, a multi-scale mask-guided encoder is designed to project the input face into the per-region style codes, which are resident in the style space of StyleGAN. Besides, we design a mask-guided injection module that uses the style codes to manipulate the feature maps in the generator according to the given masks. Besides of shape of texture, we propose to transfer the target lighting to the swapped face by training a face re-coloring network, which is trained in a self-supervised manner to recover the corresponding grayscale images. Further, as a post-processing, we design a face inpainting network to maintain the source face shape on the pixel level. We conduct extensive experiments on face swapping, face editing and some extended applications. The results and comparisons with current state-of-the-art methods demonstrate the superiority of the E4S framework, RGI method, face re-coloring and face inpainting network.

### v3

@@ III-D Training Objective
Different from the most current face swapping methods, our E4S takes reconstruction as the proxy task, making it easy to train. When training is finished, one can employ the texture encoder $F_{\phi}$ to generate per-region texture codes for any input face. Then, face swapping can be easily fulfilled as described in Sec. III-A. We apply the commonly used loss functions in the GAN inversion methods, which are detailed in our Appendix, where the training loss of the face re-coloring network $B_{\psi}$ and face inpainting network $P_{\tau}$ are also presented.
IV Experiment Setup
IV-A DataSets.
FaceForensics++ [55] is a forensics dataset consisting of 1,000 original video sequences containing 1,000 human identities. The frames are mostly frontal faces without occlusions.
@@ IV Experiment Setup
FFHQ [50] contains 70K high-quality at $1024^{2}$ resolution. It includes vast variation in terms of age, ethnicity and image background, and also has better coverage of accessories such as eyeglasses, sunglasses, hats, etc. Since it does not provide the facial segmentation masks, we use a pre-trained face parser [51] to obtain the facial segmentation masks.
IV-B Implementation Details.
The proposed RGI is trained with PyTorch [56] with 8 NVIDIA Tesla A100 GPUs. The batch size is 2 at each GPU during training. The learning rate is initialized as $10^{-4}$ with the Adam [57] optimizer ($\beta_{1}=0.9$, $\beta_{2}=0.999$). We train the model for 200K and 300K iterations on CelebAMask-HQ and FFHQ datasets, respectively. Besides, the initial learning rate decays by the factor of $0.1$ at 100K and 150K iterations, separately. Here, all images are randomly flipped with a ratio of $0.5$. As for the adversarial training, we update the parameters of the discriminator once the generator gets updated 15 times. The reason is that the pre-trained StyleGAN can well preserve the texture details of inner face components, while sometimes its performance is unstable on preserving the hair details. Hence, we fine-tune the first $K=13$ layers to improve the hair quality. As for the RGI-Optimization, we fix the learning rate as $1e^{-2}$. Empirically, approximately 50 steps of optimization would bring satisfying results.
@@
[…]
@@
The Role of Re-enactment. We first study the roles of the re-enactment step of our E4S framework, which aims to drive the source to show a similar pose and expression as the target. However, aside from pose and expression, an ideal reenactment model should not affect other source attributes, such as identity. Specifically, we employ a pre-trained face reenactment model [15] before the shape and texture swapping procedure. To verify the necessity of the re-enactment step in E4S, we compare a standard E4S pipeline and the one without re-enactment. As shown in the 1st column of Fig. 7, the swapped result is not aligned with the target face when the reenactment is disabled.
Figure 8: Qualitative results of ablation study on face re-coloring. $M_{\rm{rec}}$ denotes the low-pass mask used for re-coloring. Comparing to the 3rd column ($\tilde{I}$), the results in the 4th column ($\tilde{I}_{\rm{rec}}$) have better lighting consistency with the target while preserving the skin texture of the source, demonstrating the efficacy of our re-coloring network. However, the previous methods DiffSwap [41] in the 5th column and BlendFace [3] in the 6th column fail to preserve the source identity although they keep the target lighting.
Figure 9: Qualitative results of ablation study on face inpainting (please zoom in for more details). The green box indicates the face shape is consistent with the source image, while the red box means the face shape is inconsistent with the source. the inpainting segmentation map $M_{\rm{inp}}$ and the orange pixels in $M_{\rm{inp}}$ denotes the mismatch regions $M_{\rm{r}}$.
@@
Pre-trained vs fine-tuned StyleGAN. The pre-trained StyleGAN can be used for face swapping. However, we notice that the hair texture details cannot be always well preserved. For a more robust performance on hair, we fine-tune the first $K=13$ layers of the StyleGAN. In Table I, we use a configuration (A) to indicate freezing the parameters of the StyleGAN generator and only training the texture encoder $F_{\phi}$ and the subsequent MLPs in our RGI. Although slightly better SSIM, PSNR, and RMSE can be achieved by (A), the FID is poor. The last column in Fig. 7 also supports this. In contrast, fine-tuning can improve hair quality while maintaining the texture of other inner facial components. The generated images look unrealistic and lack certain textures (especially in hair). The generation quality can also be proved by the higher FID.
The Role of face re-coloring network $B_{\psi}$. To evaluate the proposed face re-coloring network $B_{\psi}$, we provide the qualitative results in Fig. 8. The comparison results demonstrate that using re-coloring network can produce more natural and target-consistent lighting. Besides, although $B_{\psi}$ only generates $256^{2}$-resolution images, the final re-coloring results have the comparable image quality with the $1024^{2}$ sized inputs, with the help of face enhancing network [54] restoring the details and re-coloring mask $M_{\rm{rec}}$ preserving the high-frequency information from the high-resolution inputs.
The Role of face inpainting network $P_{\tau}$. We also study the role of face inpainting network $P_{\tau}$ in Fig. 9. For clarity, we merge the mismatch region $M_{\rm{r}}$ and swapping segmentation $M_{\rm{swap}}$ into an inpainting map $M_{\rm{inp}}$ and visualize it in the last column, whose orange pixels indicate the mismatch regions. The comparison between the 2nd and 3rd column clearly shows that the proposed network $P_{\tau}$ can successfully and adaptively keep the face shape of the source, even when the face-shape difference between source and target is extremely large. As shown in the 1st row, when there are no mismatch regions (no orange pixels in $M_{\rm{inp}}$), both our methods with or without inpainting network produce the source shape consistent results. However, the state-of-the-art methods DiffSwap [41] and BlendFace [3] fail to achieve this, whose swapped faces only inherit the face shape of the target. The results in the 2nd and 3rd rows demonstrate that when the mismatch regions (orange pixels in $M_{\rm{inp}}$) exist, our inpainting network can restore the mismatched parts, leading to more consistent face shape with the driven sources (please zoom in for more details).
Besides, the quantitative comparison between our standard E4S and the one without face re-coloring network $B_{\psi}$ or face inpainting network $P_{\tau}$ are demonstrated in Table II. We will elaborate on this in Sec. VI-B.
@@
Figure 11: Quantitative evaluation on FaceForensics++ [55]. The 1st and 2nd figures demonstrate that our method achieves the best trade-off between the identity similarity and target attributes preservation. In the 3rd and 4th figures, the gray circles indicate that BlendFace [3] uses the testing model to train the network as they stated in [3]. Comparing with the rest methods, our method achieve the best trade-off in the 3rd figure and the best performance in the 4th figure.
Figure 12: Quantitative evaluation on CelebAMask-HQ [14], where the grey circle indicates that BlendFace [3] uses the testing model to train the network, which is unfair on the BlendFace Similarity [3] metric. The left three figures show that our method can achieve the best trade-off between the FID and other metrics, which demonstrates our ability to natural and consistent swapped results. The 4-th figure shows the identity similarity and the model parameters, where our method outperforms the existing methods no matter whether it is with or without the recoloring and inpainting network.
VI Face Swapping Results.
@@
We illustrate several face swapping examples synthesized by our E4S framework and several representative methods in Fig. 10. Our method can realize high-quality swapping at the resolution of 1024$\times$1024. Then, we use the 1,000 FaceForensics++ testing pairs provided by [5] and randomly sample 30,000 source-target pairs from the test set of CelebAMask-HQ to obtain the quantitative results of different approaches. Besides, we provide video face swapping results in Appendix. Moreover, we develop a user-interface system to perform image and video face swapping, which are also detailed in Appendix.
VI-A Qualitative comparison.
@@
We provide the qualitative comparisons with state-of-the-art methods in Fig. 10, where our E4S achieves more realistic and high-fidelity swapped results than the others. The results of SimSwap [4] and HifiFace [6] suffer from some artifacts and distortions (see the 2nd and 4th row). Although both our E4S and FaceShifter [5] can generate visually-satisfying results, ours shows better detailed textures. InfoSwap [58] fails to transfer the gender in the 1st and 4th row.
We further compare the performance in more challenging cases where the occlusion exists in the source and target faces (see the 3rd and 4th row in Fig. 10). It can be observed that our E4S can fill out the missing skin in the source face (see the 4nd row), and preserve the target glasses (see the 3rd row). As contrast, MegaFS [1], StyleFusion [2], BlendFace [3] and DiffSwap [41] results have many artifacts, hard to preserving the occlusions. The results of MegaFS [1] look to be a mixture of the source and target, which are blurry and lack of textures. StyleFuison [2] shows a bit of over-smoothing (see the last row). BlendFace [3] and DiffSwap [41] both show very unnatural skin texture and inconsistent source identities. As contrast, our E4S can generate more realistic and high-quality face swapping results. The results of MegaFS [1] look to be a mixture of the source and target, which are blurry and lack of textures. StyleFuison [2] shows a bit of over-smoothing (see the last row). As contrast, our E4S can generate more realistic and high-quality face swapping results.
@@
Figure 13: Qualitative comparisons with state-of-the-art face editing methods. Some modifications are made on the face contour, hair, nose, and eyebrows. Our method can produce more high-fidelity editing results while maintaining the details of other components and the overall identity information well.
TABLE II: Quantitative comparison with the existing methods on CelebAMask-HQ testing dataset [14]. We denote the best results with bold marks, the 2nd and 3rd highest scores are noted with $\underline{\rm{underline}}$ and $\dotuline{\rm{dot}}$ respectively. Besides, the gray numbers mean that the method uses the testing model to train the network, resulting in high BlendFace-related scores [3].
[…]
Top-1(%) Top-5(%) Sim. Top-1(%) Sim.
FSGAN [13] 17.00 32.18 0.2691 63.40 0.4933 2.332 3.229 30.86 $256^{2}$ 266.72
FaceShifter [5] 6.35 31.77 0.3161 48.95 0.4170 1.731 2.090 26.03 $256^{2}$ 249.50
SimSwap [4] 11.73 26.09 0.3665 $\underline{75.68}$ 0.5164 2.892 2.129 22.41 $256^{2}$ 120.21
InfoSwap [58] 20.45 50.20 0.3284 47.56 0.4124 $\underline{2.218}$ 1.512 7.34 $512^{2}$ 250.56
MegaFS [1] 29.41 45.48 0.2588 31.59 0.3286 3.044 2.162 21.49 $1024^{2}$ 338.60
HifiFace [6] 14.71 36.94 0.3656 81.70 $\underline{0.5509}$ 2.768 1.904 15.86 $256^{2}$ 258.99
StyleFusion [2] 35.09 48.76 0.3206 16.89 0.3151 6.053 1.965 38.01 $1024^{2}$ 214.89
DiffSwap [41] 3.37 24.61 0.1947 30.29 0.3341 3.183 3.766 5.56 $256^{2}$ 530.57
BlendFace [3] 9.23 35.84 0.2784 91.03 0.6525 2.976 1.573 $\underline{6.90}$ $256^{2}$ 765.40
E4S (w/o $B_{\psi}$) $\underline{40.10}$ $\underline{57.24}$ $\underline{0.3716}$ 73.10 0.5043 3.287 0.959 10.17 $1024^{2}$ 334.95
E4S (w/o $P_{\tau}$) 37.78 56.98 0.3684 74.55 0.4982 3.275 $\underline{0.910}$ 9.04 $1024^{2}$ 366.13
E4S (Ours) 42.42 58.22 0.3829 75.02 0.5684 3.281 0.881 9.02 $1024^{2}$ 423.94
VI-B Quantitative comparison.
@@
The quantitative comparison is not only illustrated in Fig. 11 and Fig. 12 to clearly show the tradeoff between different metrics, but also detailed in Table II. Note that in Fig. 11, we empirically found that the 10,000 testing pairs of FaceForensics++ [55] having 10 images per identities can lead to pretty high scores (usually over 99%), making it hard to distinguish the performance of the existing methods. Therefore, we chose to use only one image per identity, which is a harder testing protocol, leading to the lower but more distinguished scores. The evaluation protocol mainly considers two aspects: identity preservation from the source and the attribute preservation from the target. Specifically, inspired by FaceShifter [5], our experimental settings are as follows. For source identity preservation, we first extract the ID feature vectors of all the source faces and the swapped results via CosFace [59]. Then, we conduct face retrieval by searching for the most similar face from all the source images for each swapped face, where the similarity is computed as the cosine distance. We use Top-1 and Top-5 CosFace [59] retrieval accuracy, cosine similarity and Top-1 BlendFace [3] retrieval accuracy and consine similarity as the source identity consistency metrics. For the target attribute preservation, we estimate the pose and expression by using HopeNet [60] and a 3D face reconstruction model [11], respectively. The $\ell_{2}$ distance of the pose and expression between each swapped face and its ground-truth target face is used as metrics. We also calculate the FID [61] score to compare the image quality of the methods.
@@
Table II shows our E4S achieves the best retrieval accuracy, which demonstrates the superiority of our method on identity preservation. As for the target attribute preservation, our E4S achieves comparable results in pose and expression. In general, the target-oriented methods [5, 4, 58, 3] usually keep more accurate pose and expression, since they start from the target face while our E4S is a source-oriented method, which generates the target attributes starting from the source. Nonetheless, one side effect of the target-oriented methods is that the injected identity information from the source is always limited (see FaceShifter and BlendFace in Fig. 10). To sum up, there is a trade-off between identity and attribute preservation in source- and target-oriented methods. Note that our E4S has the potential to achieve better results by leveraging a more advanced reenactment method.
@@
Besides, we also evaluate how the proposed face re-coloring network $B_{\psi}$ and face inpainting network $P_{\tau}$ make a difference to the overall E4S framework in Table II. Although the face re-coloring network $B_{\psi}$ basically does not affect the performance on source ID retrieval and target pose and expression, it totally changes the visual effects of human eyes since the face skin tone is maintained from the target rather the source (see Fig. 8). Besides, Fig. 9 has demonstrated our face inpainting network $P_{\tau}$ can successfully maintain the face shape no matter the mismatch regions exist (2nd and 3rd rows) or not (1st row), which cannot be solved well by mask-swapping only. The quantitative comparison in Table II demonstrates our face inpainting network $P_{\tau}$ can improve the performance of ID retrieval since accurately maintaining source face shape is important for source ID retrieval. Such a performance improvement is consistent with the results in Fig. 9.
VII Face Editing Results Via RGI.
@@
Qualitative comparison. We show the visual comparison with the competing methods in Fig. 13. We make some modifications to the original facial mask, such as hair, eyebrows, and face contour. It shows that our approach produces more high-fidelity editing results, where the details of other components and the overall identity information are well kept.
TABLE III: Quantitative comparison for image reconstruction on CelebAMask-HQ [14] test set. The rows in gray indicate the reconstructions are obtained via style code optimization.
[…]
@@
Quantitative comparison. We make comparison between the competing methods and our RGI on the image reconstruction quality, where the Structural Similarity Index (SSIM) [63], Root-Mean-Squared-Error (RMSE), Peak Signal-to-Noise Ratio (PSNR), and Fréchet Inception Distance (FID) [61] are used as the metrics. The results are showed in Table III. We also compare E4S with SofGAN [48], which is a StyleGAN-based method and takes style code optimization for the reconstruction. For a fair comparison, we use RGI-Optimization. Table III shows that our method always achieves the best performance on all metrics, indicating the superiority of our method. SEAN [47] sometimes produces artifacts on hair regions while our RGI can achieve more high-fidelity reconstructions with better identity, texture, and illumination. Besides, our RGI-Optimization can preserve the facial details better (e.g., the curly degree of hair, the thickness of the beard, dimples, and background).
VIII Conclusion
In this paper, we present a novel framework E4S for face swapping, which considers conducting face swapping from the perspective of fine-grained face editing, i.e., “editing for swapping”. Our E4S proposes to explicitly disentangle the shape and texture of each facial component, thus the face swapping task can be reformulated as a simplified problem of texture and shape swapping. To achieve such disentanglement as well as high resolution and high fidelity, we propose a novel Regional GAN inversion (RGI) method. Specifically, a multi-scale mask-guided encoder is designed to project the input face into the per-region style codes, which are resident in the style space of StyleGAN. Besides, we design a mask-guided injection module that uses the style codes to manipulate the feature maps in the generator according to the given masks. Besides of shape of texture, we propose to transfer the target lighting to the swapped face by training a face re-coloring network, which is trained in a self-supervised manner to recover the corresponding grayscale images. Further, we design a face inpainting network to maintain the source face shape on the pixel level. We conduct extensive experiments on face swapping, face editing and some extended applications. The results and comparisons with current state-of-the-art methods demonstrate the superiority of the E4S framework, RGI method, face re-coloring and face inpainting network.